Balance the following equations, using the oxidization number method for the redox part of the equation. Show your work.


A. Cu2O(s) + H2(g) —> Cu(s) + H2O(l)


B. Cl2(g) + KBr(aq) —> Br2(l) + KCl(aq)


1
Expert's answer
2022-05-30T17:01:12-0400

A.

we write the degree of oxidation on the top each element in the reaction equation:

Cu2+O2– + H20 = Cu0 + H2+O2–

we write the pre- and post-reaction states of elements with changed oxidation states and the number of electrons involved in its change:

Cu+ + 1 e- —> Cu0 1

H20 – 2 e- —> 2H+ 2

We write the number of electrons involved in the change of the first element as a coefficient for the second element. We write the number of electrons involved in the change of the second element as the coefficient of the first element:

Cu2O + H2 = 2 Cu + H2O


B.

we write the degree of oxidation on the top each element in the reaction equation:

Cl20 + K+Br = Br20 + K+Cl

we write the pre- and post-reaction states of elements with changed oxidation states and the number of electrons involved in its change:

Cl20 + 2 e- —> 2Cl 2

2 Br – 2 e- —> Br20 2

We write the number of electrons involved in the change of the first element as a coefficient for the second element. We write the number of electrons involved in the change of the second element as the coefficient of the first element:

Cl2 + 2 KBr = Br2 + 2 KCl